A whale swimming against an ocean current traveled 60 mi in 2h.Swimming in the opposite direction,with the current,the whale was able to travel the same distance in 1.5 h.Find the speed of the whale in calm water and the rate of the ocean current.
Answer by TimothyLamb(4379)

w - c = 60 / 2 = 30 mph
w = 30 + c
w + c = 60 / 1.5 = 40 mph
---
w + c = 40
30 + c + c = 40
30 + 2c = 40
2c = 10
---
c = 5
w = 35
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php